# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v1

@@ D.1.1 Problem Statement
Fiber-radius ($r_{f}$, nm): Radius of the actual fiber/wire that forms the helix structure. The values for this parameter range from 20 nm to 60 nm, with 10 evenly spaced values.
Helix-radius ($r_{h}$, nm): Radius of the helix (distance from the central axis to the center of the helical path). The values for this parameter range from 20 nm to 90 nm, with 10 evenly spaced values.
Number of turns ($n_{t}$, dimensionless): Number of complete turns in the helix. The values for this parameter range from 3 to 10, with 10 evenly spaced values.
Pitch ($p$, nm): Axial distance between adjacent turns. The values for this parameter range from 60 nm to 200 nm, with 10 evenly spaced values.
@@
D.1.2 Development
@@
D.2 Bio-activity Optimization
D.2.1 Problem Statement
@@ D.2.2 Development
D.3 Superconductor Critical Temperature Optimization
D.3.1 Problem Statement
Superconductivity is a quantum mechanical phenomenon where certain materials exhibit zero electrical resistance and expel magnetic fields when cooled below a critical temperature ($T_{c}$) [17]. Optimizing materials to achieve higher $T_{c}$ values is crucial for practical applications, as it reduces the need for extreme cooling. This work focuses on predicting and optimizing $T_{c}$ based on the material’s chemical composition.
@@ D.3.1 Problem Statement
D.3.2 Development
Figure 8: The $R^{2}$ of the surrogate model for critical temperature prediction. The model achieves an $R^{2}$ of 0.9133.
The dataset used for training and testing the surrogate model comprises 26,321 records of superconducting materials, including their chemical formulas and experimentally determined $T_{c}$ values [17]. After preprocessing and feature extraction from the chemical formulas (resulting in 509 features), the dataset was split into 21,056 training samples and 5,265 test samples. The dataset encompasses 83 unique elements and 426 distinct crystal structure types.
@@ D.3.2 Development
Appendix E Prompt Engineering
E.1 Planner Agent
@@ 6. Suggest all valuable insights to Hypothesis Agent
E.2 Hypothesis Agent
@@
E.3 Experiment Agent

### v2

@@ 3 Explore ChemToolAgent generated a principle $p_{3}$ and proposed an invalid formula null
Fiber-radius ($r_{f}$, nm): Radius of the actual fiber/wire that forms the helix structure. The values for this parameter range from 20 nm to 60 nm.
Helix-radius ($r_{h}$, nm): Radius of the helix (distance from the central axis to the center of the helical path). The values for this parameter range from 20 nm to 90 nm.
Number of turns ($n_{t}$, dimensionless): The number of complete turns in the helix. The values for this parameter range from 3 to 10.
Pitch ($p$, nm): Axial distance between adjacent turns. The values for this parameter range from 60 nm to 200 nm.
@@
S.1.2 Objective and Benchmark
Objective. Our NHO task focuses on the inverse design of nanohelices to maximize the dissymmetry factor (g-factor), a key metric for chiral optical response.
The challenge of complex & high-dimensional design space. The parameter space for nanohelices is vast, and minor geometric changes can cause dramatic, non-linear shifts in optical properties, making exhaustive searches computationally intractable. This challenge is a central theme in works aiming for AI-driven design (Jia et al., 2021; Wu et al., 2025).
Performance benchmark. Simulation-based analysis by Wu et al. (2025) identified nanohelices with g-factors approaching 1.8. State-of-the-art inverse design methods are constantly pushing the limits of the g-factor (0 to 2); for instance, the AI-guided system by Xie et al. (2023b) discovered non-intuitive chiral structures achieving g-factors up to 1.9 in a different material system.
PiFlow’s performance in context. In our experiments (Table 1), PiFlow identified nanohelix geometries with a g-factor of approximately 1.6. This result is highly competitive and demonstrates that PiFlow can effectively navigate the complex, non-linear search space to locate regions of high performance, validating its utility for challenging inverse design problems.
S.1.3 Development
Figure 12: The $r^{2}$ of the surrogate model for g-factor prediction.
@@
S.2 Task 2: Bio-activity Optimization (MBO)
S.2.1 Problem Statement
@@
S.2.3 Objective and Benchmark
Objective. Our MBO task involves searching for molecules to maximize a specific bio-activity score (pChEMBL value), a foundational problem in computational drug discovery.
The challenge of vast chemical space and data sparsity. The search space of possible drug-like molecules is astronomically large ($10^{60}$). Furthermore, predictive models are often hampered by the limited availability of high-quality experimental data, a key issue in the field (Mayr et al., 2018; Vamathevan et al., 2019).
Performance benchmark. Performance is measured by the ability to identify potent compounds. A pChEMBL value of 6.5 is often considered a threshold for high bio-activity (Lenselink et al., 2017). While the maximum recorded value is 11.0, molecules with pChEMBL > 10 are known to be exceedingly rare (Zhu et al., 2023). Seminal works like Zhavoronkov et al. (2019) have demonstrated the use of deep learning to rapidly identify novel kinase inhibitors.
PiFlow’s performance in context. Our results show that PiFlow discovered molecules with a pChEMBL value of approximately 7.24 (Table 1). This confirms that the framework can successfully search a vast chemical space to identify novel molecules with significant biological activity, providing a strong starting point for further optimization.
S.3 Task 3: Superconductor Critical Temperature Optimization (SPO)
S.3.1 Problem Statement
Superconductivity is a quantum mechanical phenomenon where certain materials exhibit zero electrical resistance and expel magnetic fields when cooled below a critical temperature ($T_{c}$) (Hamidieh, 2018). Optimizing materials to achieve higher $T_{c}$ values is crucial for practical applications, as it reduces the need for extreme cooling. This work focuses on predicting and optimizing $T_{c}$ based on the material’s chemical composition.
@@
S.3.2 Development
Figure 14: The $r^{2}$ of the surrogate model for $T_{c}$ prediction.
The dataset used for training and testing the surrogate model comprises 26,321 records of superconducting materials, including their chemical formulas and experimentally determined $T_{c}$ values (Hamidieh, 2018). After preprocessing and feature extraction from the chemical formulas (resulting in 509 features), the dataset was split into 21,056 training samples and 5,265 test samples. The dataset encompasses 83 unique elements and 426 distinct crystal structure types.
@@
S.3.3 Objective and Benchmark
Objective. Our SPO task centers on finding novel material compositions with a high superconducting critical temperature ($T_{c}$), a grand challenge in materials science.
The challenge of combinatorial complexity & lack of guiding theory. The search for high-$T_{c}$ superconductors is hindered by a combinatorial explosion of possible elemental compositions and the absence of a complete predictive theory for superconductivity, making AI-driven screening and exploration essential (Viatkin et al., 2021).
Performance benchmark. Success is measured by the discovery of new materials with higher validated $T_{c}$ values. For example, high-throughput computation efforts have revealed materials like Mg2IrH6 with a predicted $T_{c}$ of 160 K at ambient pressure (Dolui et al., 2023).
PiFlow’s performance in context. Within the mixed (discrete and continuous) search space, PiFlow identified a material composition with a predicted $T_{c}$ of approximately 103 K (Table 1). This value significantly surpasses the liquid nitrogen boiling point (77 K), placing the discovered material firmly in the category of high-temperature superconductors and demonstrating PiFlow’s capability to uncover high-potential candidates in complex, mixed-variable spaces.
Takeaway: This work validates the PiFlow framework’s effectiveness and versatility across three distinct and challenging scientific inverse design tasks. By integrating high-fidelity surrogate models ($R^{2}>0.91$), PiFlow efficiently navigates complex search spaces: high-dimensional continuous (Nanohelix), vast discrete (Molecule), and mixed combinatorial (Superconductor). It successfully identifies high-performance candidates in each domain: 1. the nanohelix with a g-factor of $\approx 1.6$; 2. the molecule with a pChEMBL of $\approx 7.24$; 3. the superconductor with a critical temperature ($T_{c}$) of $\approx 103$ K, demonstrating its capability to accelerate discovery in diverse scientific fields.
Appendix T Agent Prompts
T.1 Planner Agent
@@ 6. Suggest all valuable insights to Hypothesis Agent
Planner Agent. The Planner Agent serves as the strategic nexus of the system. Its core function is to translate high-level insights from the PiFlow framework into actionable guidance for the Hypothesis Agent. To ensure its directives are logical and well-grounded, its behavior is constrained by a required structured output format, compelling it to synthesize historical data and articulate a clear, principle-driven research direction in each cycle.
T.2 Hypothesis Agent
@@
Rationale Design. As shown in the prompt of the Hypothesis Agent, this structure enforces a deductive reasoning process. The prompt for $\mathcal{A}_{H}$ is engineered to explicitly request these two components before stating the final hypothesis. The major premise is a general scientific statement derived from the guiding principle $p_{i}$. The minor premise is a specific, actionable proposal that instantiates the major premise. For instance, in the search for high-temperature superconductors:
  • Principle: Introducing specific dopants can alter electron-phonon coupling and increase $T_{c}$.
  • Major Premise: “Elements with a different atomic radius can create lattice strain, which is a known mechanism to influence a material’s critical temperature ($T_{c}$).”
  • Minor Premise: “Strontium (Sr) has a different atomic radius than Barium (Ba). Let’s substitute 5% of Ba with Sr in the YBa2Cu3O7 compound.”
The final hypothesis, that this specific substitution will increase $T_{c}$, is then the direct, testable conclusion. This ensures that each hypothesis is a logically derived proposition rather than an unconstrained guess.
T.3 Experiment Agent
@@ 4. Maintain a record of prior experimental outcomes
Experiment Agent ($\mathcal{A}_{E}$). The Experiment Agent acts as a dedicated executor, whose role is strictly confined to validating the hypothesis $h_{t}$ proposed by $\mathcal{A}_{H}$. Following its operational directive, $\mathcal{A}_{E}$ interfaces with the computational tool ($f^{*}(\cdot)$) to run the specified experiment and reports the quantitative outcome objectively. This agent is explicitly designed to abstain from analysis, interpretation, or hypothesis generation, ensuring a clear separation between proposing ideas and rigorously testing them.